# The local moduli of Sasaki-Einstein $\mathbb{Q}$-homology 7-spheres and invertible polynomials

### v1

@@ 2.1. Sasakian structures on links and the Berglund-Hübsch rule
where $S_{\mathbf{w}}^{2n+1}$ denotes the unit sphere with a weighted Sasakian structure, $\mathbb{P}(\mathbf{w})$ is a weighted projective space coming from the quotient of $S_{\mathbf{w}}^{2n+1}$ by a weighted circle action generated from the weighted Sasakian structure. The top horizontal arrow is a Sasakian embedding and the bottom arrow is a Kählerian embedding and the vertical arrows are orbifold Riemannian submersions.
It follows from the orbifold adjunction formula that the link $L_{f}$ admits a positive Ricci curvature if the quotient orbifold $Z_{f}$ by the natural action $S^{1}$ is Fano, which is equivalent to
$I=|{\bf w}|-d_{f}>0.$
Here $|\mathbf{w}|=\sum_{i=0}^{m}w_{i}$ denotes the norm of the weight vector $\mathbf{w}$ and $d_{f}$ is the degree of the polynomial $f$, one refers to its difference $I$ as the index.
In [22], Kobayashi showed that the link of a cone over a smooth projective variety $Z\subset\mathbb{P}^{n}$ carries a natural Einstein metric if and only if $Z$ is Fano and $Z$ carries a Kähler-Einstein metric. In [4], the authors generalized this result to weighted cones and furthermore gave an algorithm, the Kobayashi-Boyer-Galicki method, to obtain $(n-1)$-connected Sasaki-Einstein $(2n+1)$-manifolds from the existence of orbifold Fano Kähler-Einstein hypersurfaces $Z_{f}$ in weighted projective $2n$-space $\mathbb{P}(\mathbf{w}).$
In order to obtain topological information of the link, an important device is given by the Alexander polynomial $\Delta_{f}(t)$ in [27] associated to a link $L_{f}$ of dimension $12n-$ 1 is the characteristic polynomial of the monodromy map $h_{*}:H_{n}(F,\mathbb{Z})\rightarrow H_{n}(F,\mathbb{Z})$ induced by the circle action on the Milnor fibre $F$. Then $\Delta_{f}(t)=\operatorname{det}\left(t{\mathbb{I}}-h_{*}\right)$. Now both $F$ and its closure $\bar{F}$ are homotopy equivalent to a bouquet of $n$-spheres $S^{n}\vee\cdots\vee S^{n},$ and the boundary of $\bar{F}$ is the link $L_{f}$, which is $(n-2)$-connected. The Betti numbers $b_{n-1}\left(L_{f}\right)=b_{n}\left(L_{f}\right)$ equal the number of factors of $(t-1)$ in $\Delta_{f}(t)$. From the Wang sequence of the Milnor fibration (see [32])
@@
one obtain that $L_{f}$ is a $\mathbb{Q}$-homology sphere if and only if $\Delta_{f}(1)\neq 0$ and the order of $H_{n-1}\left(L_{f},\mathbb{Z}\right)$ equals $|\Delta_{f}(1)|.$ In the case that $f$ is a weighted homogeneous polynomial there is an algorithm due to Milnor and Orlik [28] to calculate the free part of $H_{n-1}\left(L_{f},\mathbb{Z}\right).$ The authors associate to any monic polynomial $f$ with roots $\alpha_{1},\ldots,\alpha_{k}\in\mathbb{C}^{*}$ its divisor
@@
Invertible polynomials and the Berglund-Hübsch transpose: In [13, 14] we used the Kobayashi-Boyer-Galicki method to establish the existence of Sasaki-Einstein metrics on links of invertible polynomials. Recall that a polynomial $f\in\mathbb{C}\left[x_{1},\ldots,x_{n}\right]$ is invertible if it is of the form
$f=\sum_{i=1}^{n}\prod_{j=1}^{n}x_{j}^{a_{ij}}$
where $A=\left(a_{ij}\right)_{i,j=1}^{n}$ is a non-negative integer-valued matrix satisfying:
  1. (1)
    $A$ is invertible over $\mathbb{Q}$;
  2. (2)
    $f$ is quasi-homogeneous, i.e., there exists positive integers $w_{j}$ such that $d:=\sum_{j=1}^{n}w_{j}a_{ij}$ is constant for all $i$; and
  3. (3)
    $f$ is quasi-smooth, i.e., $f:\mathbb{C}\left[x_{1},\ldots,x_{n}\right]\rightarrow\mathbb{C}$ has exactly one critical point at the origin.
Due to the Kreuzer-Skarke classification of invertible polynomials [25] we know that any invertible polynomial, up to permutation of variables, can be written as a Thom-Sebastiani sum of three types of polynomials usually called atoms:
  1. (1)
    Fermat type: $w=x^{r}$,
  2. (2)
  3. (3)
For instance Brieskorn-Pham polynomials $f=x_{1}^{a_{1}}+x_{2}^{a_{2}}+\ldots+x_{n}^{a_{n}}$ are Thom-Sebastiani sums of Fermat type polynomials.
In [5], [13] all the invertible polynomials that produce $\mathbb{Q}$-homology 7-spheres and admit Sasaki-Einstein metrics are polynomials of cycle type, chain type, or iterated Thom-Sebastiani sums of these types. These polynomials were taken from the list of Johnson and Kollár of anticanonically embedded Fano 3-folds of index 1 [19]. All these polynomials satisfy one of the following conditions:
@@
  • The weights are subject to $(w_{0},w_{1},w_{2},w_{3},w_{4})=(m_{2}v_{0},m_{2}v_{1},m_{2}v_{2},m_{3}v_{3},m_{3}v_{4})$ with $\gcd(m_{2},m_{3})=1$ and $m_{2}m_{3}=d,$ which leads to singularities that can described as iterated Thom-Sebastiani sums of chain, cycle type and Fermat singularities. More precisely, the types of singularities obtained have the following form:
@@
Proofs of the following statements can be found in [5] and/or in the proof of Theorem 3.2 in [13]:
@@
    Consider the link $L(\textbf{w},d)$ of a weighted homogeneous polynomial $f$ with $\operatorname{gcd}(d,w_{i})=1$ for all $i=0,\dots 4,$ then the Milnor number $m(L_{f})$ for $L_{f}$ is given by
@@
    So in the case the link is a $\mathbb{Q}$-homology 7-sphere we obtain
@@
    If the link $L(\textbf{w},d)$ is such that its weight vector $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ satisfies $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})=(m_{2}v_{0},m_{2}v_{1},m_{2}v_{2},m_{3}v_{3},m_{3}v_{4})$, $\gcd(m_{2},m_{3})=1$ and $m_{2}m_{3}=d.$ One obtains the equality
@@
    $\alpha(\textbf{w})=\frac{m_{2}}{v_{3}v_{4}}-\frac{1}{v_{3}}-\frac{1}{v_{4}}$ (2.1.2)
@@
    $\beta(\textbf{w})=\left(\frac{m_{3}}{v_{0}v_{1}}-\frac{1}{v_{1}}-\frac{1}{v_{0}}\right)\left(\frac{m_{3}}{v_{2}}-1\right)+\frac{1}{v_{2}}.$ (2.1.3)
    It is known that if the link is a $\mathbb{Q}$-homology sphere, then $\beta({\bf w})=1.$
@@
    So if the corresponding link is a $\mathbb{Q}$-homology sphere, it follows that
@@
Recall that the Berglund-Hübsch transpose rule considers an invertible polynomial $f=\sum_{j=0}^{n}\prod_{i=0}^{n}x_{i}^{a_{ij}}$ cutting out an orbifold of degree $d$ in $\mathbb{P}(\bf w)$ and defines the transpose polynomial $f^{T}$ by transposing the exponential matrix $A=\left(a_{ij}\right)_{i,j}$ of the original polynomial, that is,
$f^{T}=\sum_{i=1}^{n}\prod_{j=1}^{n}x_{j}^{a_{ji}},$
again an invertible polynomial that cuts out an orbifold of degree $\tilde{d}$ in $\mathbb{P}(\tilde{\bf w}).$ Then one considers the links $L_{f}(\mathbf{w},d)$ and $L_{f^{T}}(\tilde{\mathbf{w}},\tilde{d})$ associated to each of these polynomials. We will sometimes say that these two links are Berglund-Hübsch duals to one another. The following diagram succinctly summarizes the procedure described above, where BH denotes the Berglund-Hübsch transpose rule:
${{f=0}}$${{f^{T}=0}}$${L_{f}(\mathbf{w},d)}$${L_{f^{T}}(\tilde{\mathbf{w}},\tilde{d}).}$$\scriptstyle{BH}$$\scriptstyle{BH}$
In [14] the Berglund-Hübsch transpose rule is used to produce links with the $\mathbb{Q}$-homology of a 7-sphere and it was found that this rule only produces twins for singularities of cycle type or of type I and type III. Recall [5, 14] that two links $L_{f}$ and $L_{g}$ of an isolated hypersurface singularity are called twins if both are $\mathbb{Q}$-homology $(2n+1)$-spheres and they satisfy $m\left(L_{f}\right)=m\left(L_{g}\right),d_{g}=d_{f}$, and $H_{n}\left(L_{f},\mathbb{Z}\right)=H_{n}\left(L_{g},\mathbb{Z}\right)$. However, for case II, the Berglund-Hübsch transpose rule does not preserve neither torsion nor Milnor number. Its dual links will receive special attention is Subsection 3.3.
@@ 2.2. Deformations of transverse holomorphic Sasakian structures
Locally the moduli space of Sasaki isotopy classes is determined by the deformation theory of the transverse holomorphic structure of the foliation $\mathcal{F}.$ So the usual thing to do is to fix the contact structure and deform the transverse holomorphic structure via Kodaira-Spencer theory.
The cohomological groups with respect to the sheaf of holomorphic vector fields $T_{X}$ (on a complex manifold $X$) have precise interpretation in terms of deformation theory [23]: $H^{1}\left(X,T_{X}\right)$ can be seen as the space of infinitesimal deformations of the complex structure of $X,H^{2}\left(X,T_{X}\right)$ is the space of obstructions to lifting an infinitesimal deformation to an actual deformation of the complex structure of $X$. The group $H^{0}\left(X,T_{X}\right)$ can be considered as the Lie algebra of the group of holomorphic automorphisms of $X$, and if $H^{0}\left(X,T_{X}\right)=0$ then the Kuranishi family is universal. We have analogs to these notions for the transverse holomorphic case, we give a brief review of these notions following [17] for the case of interest: a Sasakian manifold $(M,\xi,\eta,\Phi,g)$ with transverse holomorphic foliation $\mathcal{F}_{\xi}.$
@@
A germ of a deformation of a transverse holomorphic foliation $\mathcal{F}_{\xi}$ on $M$ with base space $(B,0)$ is given by an open cover $\left\{U_{\alpha}\right\}$ of $M$ and a family of local submersions $f_{\alpha,t}:U_{\alpha}\rightarrow\mathbb{C}^{n}$ parametrized by $(B,0)$ that are holomorphic in $t\in B$ for each $x\in U_{\alpha}$. For $\Theta_{\mathcal{F}_{\xi}}$ denoting the sheaf of transversely holomorphic vector fields on $M$, we have a Kodaira-Spencer map $\rho:T_{0}B\rightarrow H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$ that sends $\frac{\partial}{\partial t}$ to certain class in $H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$ defined by a section $\theta_{\alpha,\beta}$ of the sheaf $\Theta_{\mathcal{F}_{\xi}}\mid U_{\alpha}\cap U_{\beta}$. One can consider the full cohomology ring $H^{*}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$, these were proven to be finite dimensional.
In [17] it is shown that there is a versal Kuranishi space of deformations given by the map $\Phi:U\rightarrow H^{2}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$, for $U$ open set in $H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$, here, as in the complex case, the base of parametrizations is given by $\Phi^{-1}(0)$. We have that if $H^{2}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)=0$, then the Kuranishi family of deformations of $\mathcal{F}_{\xi}$ is isomorphic to an open set in $H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$. Otherwise, the Kuranishi space may be singular. For a quasi-regular Sasakian structure $\mathcal{S}=(\xi,\eta,\Phi,g)$ with quotient orbifold $\mathcal{Z}$ one obtains the following sequence [8]
$0\longrightarrow H^{1}\left(\mathcal{Z},\Theta_{\mathcal{Z}}\right)\longrightarrow H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)\longrightarrow H^{0}\left(\mathcal{Z},\Theta_{\mathcal{Z}}\right)\longrightarrow H^{2}\left(\mathcal{Z},\Theta_{\mathcal{Z}}\right)$
@@
where $\Theta_{\mathcal{Z}}$ denotes the sheaf of germs of holomorphic vector fields on the complex orbifold $\mathcal{Z}$. Thus the deformation of the transverse holomorphic structures of the foliation can be understood in terms of the deformation of the complex structure of the orbifold, which are described by $H^{1}\left(\mathcal{Z},\Theta_{\mathcal{Z}}\right)$ and in terms of the deformations of the Reeb vector in $H^{0}\left(\mathcal{Z},\Theta_{\mathcal{Z}}\right)$ which are described by the Sasaki cone [9]. Of course, under the absence of automorphism on the orbifold we obtain an isomorphisms between $H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$ and $H^{1}\left(\mathcal{Z},\Theta_{\mathcal{Z}}\right).$
Also, as shown in [30], deforming the transversely holomorphic structure may not stay Sasakian unless the $(0,2)$ component of the basic Euler class $\left[d\eta^{0,2}\right]\in H^{(0,2)}_{B}(M,\mathcal{F}_{\xi})$ is zero, where $H^{0,2}_{B}(M,\mathcal{F}_{\xi})$ is the basic Dolbeault cohomology for the transversal complex structure. In case the Sasakian structure is positive, which is the case we are interested in this article, it has been proven in [30] that $H^{0,q}_{B}(M,\mathcal{F}_{\xi})=0$ for $q>0.$
@@
We pay particular attention to the local moduli of quasi-regular Sasakian structures on links of isolated singularities, so our focus is on orbifolds that are quasi-smooth weighted hypersurface in certain weighted projective space. For this kind of orbifolds the group $H^{1}\left(\mathcal{Z},\Theta_{\mathcal{Z}}\right)$ has a precise description. The proof of the next two theorems can be found in [8]. (See also [7].)
@@
Let $Z_{f}$ be a quasi-smooth weighted hypersurface in $\mathbb{CP}(\mathbf{w})$ corresponding to the weighted homogenous polynomial $f$ of degree $d$ and weight vector $\mathbf{w}=\left(w_{0},\ldots,w_{n}\right)$. Let $\mathfrak{Aut}\left(Z_{f}\right)$ denote the group of automorphisms of $Z_{f}$. Assume also that $n\geqslant 3$. Then the complex orbifolds $\mathcal{Z}=\left(Z_{f},\mathcal{U}\right)$ form a continuous family of complex dimension
$h^{0}(\mathbb{CP}(\mathbf{w}),\mathcal{O}(d))-\sum_{i}h^{0}\left(\mathbb{P}(\mathbf{w}),\mathcal{O}\left(w_{i}\right)\right)+\operatorname{dim}\mathfrak{Aut}\left(Z_{f}\right).$ (2.2.1)
Furthermore, if the index $I=|\mathbf{w}|-d>0$ and $Z_{f}$ admits a Kähler-Einstein metric for a generic $f$ then it admits a $2\left[h^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(d))-\sum_{i}h^{0}\left(\mathbb{P}(\mathbf{w}),\mathcal{O}\left(w_{i}\right)\right)+\operatorname{dim}\mathfrak{Aut}\left(Z_{f}\right)\right]$ dimensional family of Kähler-Einstein metrics up to homothety.
Here $h^{0}=\dim_{\mathbb{C}}H^{0}$ denotes the complex dimension of the corresponding space of holomorphic sections and $\mathfrak{Aut}\left(Z_{f}\right)$ denotes the group of complex automorphisms of the orbifold $Z_{f}.$
For quasi-regular Sasaki- Einstein metrics we have the following
@@
Let $M$ be a smooth compact manifold and let $\mathcal{S}=\left(\xi,\eta,\Phi_{t},g_{t}\right)$ be a family of quasi-regular Sasaki-Einstein structures on $M$ induced by a continuous family of inequivalent complex orbifolds $\mathcal{Z}_{t}$ with Kähler-Einstein metrics. Then the metrics $g_{t}$ are inequivalent as Einstein metrics.

### v2

@@ 2.1. Sasakian structures on smooth links and invertible polynomials
where $S_{\mathbf{w}}^{2n+1}$ denotes the unit sphere with a weighted Sasakian structure, $\mathbb{P}(\mathbf{w})$ is a weighted projective space coming from the quotient of $S_{\mathbf{w}}^{2n+1}$ by a weighted circle action generated from the weighted Sasakian structure. The top horizontal arrow is a Sasakian embedding, the bottom arrow is a Kählerian embedding and the vertical arrows are orbifold Riemannian submersions.
It follows from the orbifold adjunction formula that the link $L_{f}$ admits a positive Ricci curvature if the quotient orbifold $Z_{f}$ by the natural $S^{1}$-action is Fano.
In [26], Kobayashi showed that the link of a cone over a smooth projective variety $Z\subset\mathbb{P}^{n}$ carries a natural Einstein metric if and only if $Z$ is Fano and $Z$ carries a Kähler-Einstein metric. In [4], the authors generalized this result to weighted cones and furthermore gave an algorithm, the Kobayashi-Boyer-Galicki method, to obtain (n1)(n-1)-connected Sasaki-Einstein $(2n+1)$-manifolds from the existence of orbifold Fano Kähler-Einstein hypersurfaces $Z_{f}$ in weighted projective $2n$-space $\mathbb{P}(\mathbf{w}).$
Important topological information of the link can be obtained via the Alexander polynomial. Recall that the Alexander polynomial $\Delta_{f}(t)$ [37] associated to a link $L_{f}$ of dimension (2n1)(2n-1) is the characteristic polynomial of the monodromy map $h_{*}:H_{n}(F,\mathbb{Z})\rightarrow H_{n}(F,\mathbb{Z})$ which is induced by the circle action on the Milnor fibre $F$. Then $\Delta_{f}(t)=\operatorname{det}\left(t{\mathbb{I}}-h_{*}\right)$. Now both $F$ and its closure $\bar{F}$ are homotopy equivalent to a bouquet of $n$-spheres $S^{n}\vee\cdots\vee S^{n},$ and the boundary of $\bar{F}$ is the link $L_{f}$, which is (n2)(n-2)-connected. The Betti numbers $b_{n-1}\left(L_{f}\right)=b_{n}\left(L_{f}\right)$ equal the number of factors of (t1)(t-1) in $\Delta_{f}(t)$. From the Wang sequence of the Milnor fibration (see [44])
@@
one obtains that $L_{f}$ is a $\mathbb{Q}$-homology sphere if and only if $\Delta_{f}(1)\neq 0$ and the order of $H_{n-1}\left(L_{f},\mathbb{Z}\right)$ equals $|\Delta_{f}(1)|.$ In the case that $f$ is a weighted homogeneous polynomial there is an algorithm due to Milnor and Orlik [38] to compute the Alexander polynomial in terms of the degree and the weights: associate to any monic polynomial $f$ with roots $\alpha_{1},\ldots,\alpha_{k}\in\mathbb{C}^{*}$ its divisor
@@
Invertible polynomials and rational homology spheres: In [14, 16] we used the Kobayashi-Boyer-Galicki method to establish the existence of Sasaki-Einstein metrics on links of invertible polynomials. From [5] and [14] one notices that all the invertible polynomials taken from the list of Johnson and Kollár of anticanonically embedded Fano 3-folds [24] producing Sasaki-Einstein rational homology 7-spheres were polynomials of cycle type, chain type, or iterated Thom-Sebastiani sums of these types. These polynomials can be described in terms of the following two sets of conditions
@@
  • The weights are subject to $\left(w_{0},w_{1},w_{2},w_{3},w_{4}\right)=\left(m_{3}v_{0},m_{3}v_{1},m_{2}v_{2},m_{2}v_{3},m_{2}v_{4}\right)$ with $\gcd(m_{2},m_{3})=1$ and $m_{2}m_{3}=d,$ which leads to singularities that can described as iterated Thom-Sebastiani sums of chain, cycle type and Fermat singularities. More precisely, the types of singularities obtained have the following form:
@@
Polynomials with this set of conditions satisfy the following statements whose proofs can be found in [5] and/or in the proof of Theorem 3.2 in [14]:
@@
    Consider links $L(\textbf{w},d)$ of weighted homogeneous polynomials $f$ of the first kind, that is, with $\operatorname{gcd}(d,w_{i})=1$ for all $i=0,\dots 4,$ then the Milnor number $m(L_{f})$ for $L_{f}$ is given by
@@
    So in the case the link is a rational homology 7-sphere we obtain
@@
    Consider links $L(\textbf{w},d)$ of the second kind, that it, such that the weight vectors $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ satisfy $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})=(m_{3}v_{0},m_{3}v_{1},m_{2}v_{2},m_{2}v_{3},m_{2}v_{4})$, $\gcd(m_{2},m_{3})=1$ and $m_{2}m_{3}=d.$ One obtains the equality
@@
    $\alpha(\textbf{w})=\frac{m_{2}}{v_{0}v_{1}}-\frac{1}{v_{0}}-\frac{1}{v_{1}}$ (2.1.2)
@@
    $\beta(\textbf{w})=\left(\frac{m_{3}}{v_{2}v_{3}}-\frac{1}{v_{3}}-\frac{1}{v_{2}}\right)\left(\frac{m_{3}}{v_{4}}-1\right)+\frac{1}{v_{4}}.$ (2.1.3)
    It is known that if the link is a rational homology sphere, then $\beta({\bf w})=1.$
@@
    Since the corresponding link is a $\mathbb{Q}$-homology sphere, it follows that
@@
The Berglund-Hübsch transpose rule: Recall that the Berglund-Hübsch transpose rule considers an invertible polynomial
$f=\sum_{i=0}^{n}\prod_{j=0}^{n}x_{j}^{a_{ij}}$
cutting out an orbifold of degree $d$ in $\mathbb{P}(\bf w)$ and defines the transpose polynomial $f^{T}$ by transposing the exponential matrix $A=\left(a_{ij}\right)$ of the original polynomial, that is,
$f^{T}=\sum_{i=0}^{n}\prod_{j=0}^{n}x_{j}^{a_{ji}},$
[…]
In [16] the Berglund-Hübsch transpose rule is used to produce Sasaki-Einstein links with the $\mathbb{Q}$-homology of a 7-sphere. In particular, we found that this rule only produces twins for singularities of cycle type or of type I and type III. Recall [5, 16] that two links $L_{f}$ and $L_{g}$ of an isolated hypersurface singularity are called twins if both are $\mathbb{Q}$-homology $(2n+1)$-spheres and they satisfy $m\left(L_{f}\right)=m\left(L_{g}\right),d_{g}=d_{f}$, and $H_{n}\left(L_{f},\mathbb{Z}\right)=H_{n}\left(L_{g},\mathbb{Z}\right)$. However, for polynomials of type II, the Berglund-Hübsch transpose rule does not preserve neither torsion nor Milnor number. Their dual links will receive special attention is Subsection 3.3.
@@ 2.2. Deformations of transverse holomorphic Sasakian structures
Locally the moduli space of Sasaki isotopy classes is determined by the deformation theory of the transverse holomorphic structure of the foliation $\mathcal{F}_{\xi}.$ So the usual thing to do is to fix the contact structure and deform the transverse holomorphic structure via Kodaira-Spencer theory.
@@
A germ of a deformation of a transverse holomorphic foliation $\mathcal{F}_{\xi}$ on $M$ with base space $(B,0)$ is given by an open cover $\left\{U_{\alpha}\right\}$ of $M$ and a family of local submersions $f_{\alpha,t}:U_{\alpha}\rightarrow\mathbb{C}^{n}$ parametrized by $(B,0)$ that are holomorphic in $t\in B$ for each $x\in U_{\alpha}$. For $\Theta_{\mathcal{F}_{\xi}}$ denoting the sheaf of transversely holomorphic vector fields on $M$, we have a Kodaira-Spencer map $\rho:T_{0}B\rightarrow H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$ that sends $\frac{\partial}{\partial t}$ to a certain class in $H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$ defined by a section $\theta_{\alpha,\beta}$ of the sheaf $\Theta_{\mathcal{F}_{\xi}}\mid U_{\alpha}\cap U_{\beta}$. One can consider the full cohomology ring $H^{*}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$, these were proven to be finite dimensional. In [20] it is shown that there is a versal Kuranishi space of deformations given by the map $\Phi:U\rightarrow H^{2}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$, for $U$ open set in $H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$, here, as in the complex case, the base of parametrizations is given by $\Phi^{-1}(0)$. We have that if $H^{2}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)=0$, then the Kuranishi family of deformations of $\mathcal{F}_{\xi}$ is isomorphic to an open set in $H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$. Otherwise, the Kuranishi space may be singular. For a quasiregular Sasakian structure $\mathcal{S}=(\xi,\eta,\Phi,g)$ with quotient orbifold Z{Z} one obtains the following sequence [9]
$0\longrightarrow H^{1}\left({Z},\Theta_{{Z}}\right)\longrightarrow H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)\longrightarrow H^{0}\left({Z},\Theta_{{Z}}\right)\longrightarrow H^{2}\left({Z},\Theta_{{Z}}\right)$
@@
where $\Theta_{{Z}}$ denotes the sheaf of germs of holomorphic vector fields on the complex orbifold Z{Z} and $H^{0}\left({Z},\Theta_{{Z}}\right)$ can be considered as the Lie algebra of the group of holomorphic automorphisms of Z.{Z}. Thus the deformation of the transverse holomorphic structures of the foliation can be understood in terms of the deformation of the complex structure of the orbifold, which are described by $H^{1}\left({Z},\Theta_{{Z}}\right)$ and in terms of the deformations of the Reeb vector in $H^{0}\left({Z},\Theta_{{Z}}\right)$ which are described by the Sasaki cone [10]. Of course, if $H^{0}\left({Z},\Theta_{{Z}}\right)=0$ we obtain an isomorphisms between $H^{1}\left(M,\Theta_{\mathcal{F}_{\xi}}\right)$ and $H^{1}\left({Z},\Theta_{{Z}}\right).$ Deformations of the transversely holomorphic structure may not remain Sasakian unless the $(0,2)$ component of the basic Euler class $\left[d\eta^{0,2}\right]\in H^{(0,2)}_{B}(M,\mathcal{F}_{\xi})$ is zero, where $H^{0,2}_{B}(M,\mathcal{F}_{\xi})$ is the basic Dolbeault cohomology for the transversal complex structure. However, in case the Sasakian structure is positive, which is the case we are interested in this article, it has been proven in [40] that $H^{0,q}_{B}(M,\mathcal{F}_{\xi})=0$ for $q>0.$
@@
We pay particular attention to the local moduli of quasiregular Sasakian structures on links of isolated singularities, so we focus on orbifolds that are quasismooth weighted hypersurfaces in certain weighted projective spaces. Moreover, all the orbifolds under discussion in this article have finite automorphism group, so we will assume $H^{0}\left({Z},\Theta_{{Z}}\right)=0.$ The proof of the next two theorems can be found in [9]. (See also [7].)
@@
Let $Z_{f}$ be a quasismooth weighted hypersurface in $\mathbb{P}(\mathbf{w})$ corresponding to the weighted homogenous polynomial $f$ of degree $d$ and weight vector $\mathbf{w}=\left(w_{0},\ldots,w_{n}\right)$ with $H^{0}\left(Z_{f},\Theta_{Z_{f}}\right)=0.$ Assume also that $n\geqslant 3$. Then the complex orbifolds $Z_{f}$ form a continuous family of complex dimension
$\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}({d}))-\sum_{i}\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i})).$ (2.2.1)
Furthermore, if the index $I=|\mathbf{w}|-d>0$ and $Z_{f}$ admits a Kähler-Einstein metric for a generic $f$ then it admits a $2\left[h^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(d))-\sum_{i}h^{0}\left(\mathbb{P}(\mathbf{w}),\mathcal{O}\left(w_{i}\right)\right)\right]$ dimensional family of Kähler-Einstein metrics up to homothety.
For quasiregular Sasaki-Einstein metrics we have the following:
@@
Let $M$ be a smooth compact manifold and let $\mathcal{S}=\left(\xi,\eta,\Phi_{t},g_{t}\right)$ be a family of quasiregular Sasaki-Einstein structures on $M$ induced by a continuous family of inequivalent complex orbifolds $\mathcal{Z}_{t}$ with Kähler-Einstein metrics. Then the metrics $g_{t}$ are inequivalent as Sasaki-Einstein metrics.